# Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis, Solution and Interpretation

### v1

@@ Abstract
Previous studies show that introducing new knowledge during large language models (LLMs) fine-tuning can lead to the generation of erroneous output when tested on known information, thereby triggering factual hallucinations. However, existing studies have not deeply investigated the specific manifestations and underlying mechanisms of these hallucinations. Our work addresses this gap by designing a controlled dataset Biography-Reasoning, and conducting a fine-grained analysis across multiple knowledge types and two task types, including knowledge question answering (QA) and knowledge reasoning tasks. We find that when fine-tuned on a dataset in which a specific knowledge type consists entirely of new knowledge, LLMs exhibit significantly increased hallucination tendencies. This suggests that the high unfamiliarity of a particular knowledge type, rather than the overall proportion of new knowledge, is a stronger driver of hallucinations, and these tendencies can even affect other knowledge types in QA tasks. To mitigate such factual hallucinations, we propose KnownPatch, which patches a small number of known knowledge samples in the later stages of training, effectively alleviating new-knowledge-induced hallucinations. Through attention analysis, we find that learning new knowledge reduces the model’s attention to key entities in the question, thus causing excessive focus on the surrounding context, which may increase the risk of hallucination. Moreover, the attention pattern can propagate to similar contexts, facilitating the spread of hallucinations to textually similar questions. Our method effectively mitigates the disruption of new knowledge learning to the model’s attention on key entities, accompanied by improved performance.
Codehttps://github.com/NJUNLP/New-Knowledge-Induced-Factual-Hallucinations
@@ 1 Introduction
LLMs embed rich factual knowledge in their parameters during pre-training on massive text corpora (Petroni et al. 2019; Cohen et al. 2023). Subsequently, post-training enables them to learn to follow human instructions and exhibit superior performance across various downstream tasks (Ouyang et al. 2022; Wei et al. 2022).
@@
Figure 1: The impact of learning new knowledge on attention patterns and hallucination behavior. When an model is trained on unknown facts, it may be more prone to produce factual hallucinations, the severity of which correlates with the degree of attention paid to key entities and is modulated by contextual similarity. By injecting a small amount of known knowledge at the end of training via KnowPatch, this issue can be effectively mitigated.
@@
However, during the Supervised Fine-Tuning (SFT) phase, models may encounter new knowledge not covered in pre-training. Prior researches (Ghosal, Hashimoto, and Raghunathan 2024; Lin et al. 2023; Ovadia et al. 2023; Gekhman et al. 2024; Sun et al. 2025) suggest that introducing new knowledge in the post-training phase increases the risk of factual hallucinations, where models generate fabricated yet plausible statements. This occurs because, when models learn new facts absent from pre-training, they may erroneously generate related information in irrelevant contexts (Gekhman et al. 2024; Sun et al. 2025). These studies primarily focus on the effects within knowledge-intensive QA tasks during SFT, and we advance this line of research by conducting a systematic analysis of fine-grained manifestations and underlying causes of hallucinations.
@@
To enable a comprehensive investigation, we construct a controlled experimental dataset, Biography-Reasoning. The dataset is composed of biographical entities and their four attributes (birth, death, major, and university), which serve as four knowledge types. We further design twelve reasoning tasks using these knowledge. By controlling the proportion of known and unknown knowledge within different types and tasks in the training data, we systematically investigate the impact of learning new knowledge on hallucination risks.
@@
Our experiments reveal that, for knowledge QA tasks, training on unknown knowledge significantly elevates hallucination risks in the same type test set, with some cross-type influence on other QA test sets; in tasks involving knowledge reasoning, training on reasoning tasks containing unknown knowledge primarily affects the same reasoning task, with some impact on knowledge QA and limited influence on other reasoning tasks. Importantly, we find that if a knowledge type consists entirely of new knowledge, even a small amount of such data can markedly increase hallucination tendencies. This poses challenges for non-perfect filtering-based hallucination mitigation methods: inadvertently retained small amounts of unknown data can trigger severe hallucinations.
To address this challenge, we propose KnownPatch, a simple yet effective strategy that places a small number of known knowledge in later stages of training. Instead of relying on exhaustive filtering of unknown content, our method stabilizes learning by reinducing known knowledge late in training. Experiments show that even a small injection of known data significantly reduces hallucination tendencies across diverse settings.
Finally, we perform an interpretability analysis of the model’s attention distribution. The results reveal that learning new knowledge significantly weakens the model’s attention to key entities in the question, instead may erroneously bind new knowledge to other contexts, thereby triggering factual hallucinations. Therefore, tasks that share similar contexts are more likely to be affected and exhibit stronger hallucination tendencies. KnownPatch effectively restores and enhances the model’s attention to key entities, thereby mitigating hallucination risks. These findings are visually presented in Figure 1.
In summary, the main contributions of this paper are:
@@
  • Mitigation Strategy: The introduction of KnownPatch, a training method that places a small amount of known knowledge in the later stages, significantly reducing hallucination risks without data modifications.
  • Mechanism Interpretability: An analysis of attention mechanisms shows that learning new knowledge reduces attention to key question entities and erroneously binds new knowledge to contexts, causing hallucinations. Moreover, similar contexts facilitate the spread of these altered attention patterns, enabling cross-task hallucination effects.
@@ 2 Related Work
Existing studies have indicated that introducing new knowledge into LLMs may trigger hallucinations (Ghosal, Hashimoto, and Raghunathan 2024; Lin et al. 2023; Ovadia et al. 2023). Subsequent works have provided deeper analyses of this phenomenon(Gekhman et al. 2024; Kang et al. 2024; Sun et al. 2025). Gekhman et al. (2024) found that as the proportion of new knowledge in fine-tuning data increases, the model’s hallucination tendency intensifies. Kang et al. (2024) analyzed that when fine-tuned LLMs encounter unknown queries during testing, their responses imitate those associated with unknown examples in the fine-tuning data. Sun et al. (2025), from the perspective of token probabilities, examined how, after learning new knowledge, the generation probability of relevant answer entity tokens in irrelevant contexts increases significantly, suggesting that the model may erroneously generalize knowledge, leading to hallucinations. However, previous studies focus mainly on closed-book QA settings with mixed knowledge types during training, while our controlled setup disentangles them to provide a more detailed analysis of new knowledge-induced hallucinations across types and tasks. Furthermore, we also investigate the underlying mechanisms of these phenomenon through an analysis of attention weights.
@@
Numerous studies are currently exploring ways to mitigate model hallucinations. A common approach involves providing additional relevant context to the model to reduce hallucinations during generation, such as through retrieval from knowledge bases or leveraging other large models to generate context(Shuster et al. 2021; Sun et al. 2022; Asai et al. 2024; Feng et al. 2023). Additionally, some research explicitly avoids hallucination risks by refusing to answer uncertain or unfamiliar questions(Yadkori et al. 2024; Zhu et al. 2025; Duwal 2025). In another direction, many studies encourage the model to generate more known knowledge from pre-training, for example, by promoting factual outputs via reinforcement learning (Rafailov et al. 2023; Kang et al. 2024; Li and Ng 2025; Gu et al. 2025) or by training only on known knowledge during supervised fine-tuning (Lin et al. 2024; Ghosal, Hashimoto, and Raghunathan 2024; Liu et al. 2024) to enhance the model. Our work builds on SFT with known knowledge approach, but rather than pursuing comprehensive filtering across all training data, KnownPatch only introduces a small number of known knowledge samples in the later stages of training, and alleviates the model’s tendency for hallucination.
@@ 3 Methodology of Analyzing Hallucinations
We aim to systematically investigate factual hallucinations in LLMs caused by learning different knowledge-related tasks. However, in real-world datasets, most factual knowledge may have already been seen by LLMs during pre-training, making it difficult to precisely control whether the knowledge being learned is new to the model. To address this limitation, we construct a controlled experimental environment and a synthetic dataset named Biography-Reasoning, which allows a controllable examination of hallucination behaviors under varying knowledge types and task types.
@@ 3.1 Biography-Reasoning Dataset
Following the data construction methodologies of Allen-Zhu and Li (2024); Zheng et al. (2025), we design the Biography-Reasoning dataset. The dataset centers on individuals as the key entities, with each person associated with four attributes: birth year (B), death year (D), major (M), and university (U). We refer to the same attribute of different individuals as a knowledge type.
Our dataset includes two main types of knowledge-related tasks, i.e. knowledge QA and knowledge-based reasoning tasks. For knowledge QA tasks, questions are formulated by directly querying one of the attributes given the person’s name. Each task consists of questions on a single type, resulting in four QA tasks (B_QA, D_QA, M_QA and U_QA).
For knowledge-based reasoning tasks, we design three fundamental types of chain-of-thought-requiring reasoning tasks. Specifically, these include:
  • SR (Single Reasoning): extracting one attribute from a single entity and performing a simple reasoning process;
  • CR (Comparative Reasoning): extracting one attribute from each of two entities and performing comparative reasoning between them;
  • NR (Novel Reasoning): extracting one attribute from a single entity and performing a newly defined reasoning task, such as mathematical or symbolic reasoning.
Table 1 presents examples of the constructed questions. Some of the reasoning tasks are intentionally designed to be more complex than mere knowledge extraction as QA problems. They require further reasoning, as well as auxiliary knowledge (e.g., the Major Dentistry belongs to the field Medicine), which the model is expected to contain. To further guarantee the model’s proficiency, we additionally collect and train on these auxiliary facts.
For each knowledge type we construct one QA and three reasoning tasks, leading to a total of four QA and 12 reasoning tasks per individual. Further dataset details can be found in Appendix A.
@@
M_QA Question: What major did Darreus Hsiao study?
@@
M_SR Question: What field does Darreus Hsiao’s major belong to?
@@ 3.2 Controlled Study Design
Since initially the model has no exposure to any knowledge of our synthetic dataset, we prepare the study by continue-pretraining the model with a subset of the knowledge, which becomes known to the model; and keep another subset of the knowledge as unknown. By mixing the constructed questions from known and unknown-knowledge in varying proportions, we are able to create situations where different proportion of newly introduced knowledge participates in training.
To evaluate how training leads to hallucinations, we reserve another subset of knowledge as test knowledge. The test knowledge are continue-pretrained together with the known knowledge during the preparation, but are kept away from further training. Therefore, the difference in performance on test set before and after training indicates the influence of factual hallucinations affected by training. In addition, we use the real-world ENTITYQUESTIONS dataset (Sciavolino et al. 2021) derived from Wikidata (Vrandečić and Krötzsch 2014) (denoted as wiki) as an out-of-distribution (OOD) test set to provide a more robust evaluation.
@@ 3.3 Models and Setups
We conduct experiments primarily using the Qwen2.5-1.5B model (Team 2024). As supplementary validation, we also perform partial experiments on Qwen3-8B (Team 2025) and Llama3.2-1B (Grattafiori et al. 2024) to assess generalization across model scales and architectures, with their results provided in Appendix G.
@@
Following Allen-Zhu and Li (2024) and Gekhman et al. (2024), we adopt Exact Match (EM) as the metric for both knowledge QA tasks and reasoning tasks to assess the accuracy of the final answers. Given that all test knowledge are known to the model, and the training and testing formats are consistent, there are no cases where the answers are correct but incorrectly formatted. We report the standard deviation of accuracy where applicable.
@@ 4 Hallucination Analysis
Using the Biography-Reasoning dataset, we conduct a systematic study on factual hallucinations induced by learning different tasks containing various types of new knowledge through SFT.
@@ 4.1 Knowledge QA
In this section, we analyze the impact of training on new knowledge in QA tasks. The baseline model is trained on samples constructed from the known knowledge of all four types. We then replace the knowledge of one entire type with unknown samples while keeping the other three types unchanged, resulting in four variant models. For each variant, we evaluate performance on three groups of QA test sets: (1) Same-Type QA (STQA): the test set corresponding to the replaced knowledge type; (2) Different-Type QA (DTQA): test sets of the remaining three types; (3) wiki: the real-world QA test set.
@@
-56.40 ± 4.28 -1.06 ± 0.30 -9.17 ± 4.17
Table 2: Average performance degradation (%, mean ± std) of four model variants, measured against the baseline model on different QA test groups. Detailed numerical results are reported in Appendix C.
Learning new knowledge induces factual hallucinations within the same type, with some spillover effects to other types. Table 2 presents the performance drop averaged across the four variant models. Training on unknown knowledge leads to substantial performance drops on the STQA test set, reducing the accuracy by more than half. We also observe cross-type degradation, as training on one type negatively impacts average performance on others, including the real-world wiki test set containing purely OOD knowledge. This confirms that learning new knowledge can induce hallucinations even on unrelated knowledge. Notably, the performance drop on DTQA is smaller than on wiki, as the former consists entirely of known data in the training set, which greatly mitigates the effect.
We further investigate how varying the proportion of unknown knowledge within a single type influences hallucination tendencies. Starting from the fully known-knowledge baseline, we progressively replace 5%, 10%, 20%, 50%, 80%, and 100% of the knowledge in one type with its corresponding unknown knowledge, while still keeping the other three types entirely known. Two strategies are considered for handling the remaining known knowledge within the modified type: KeepKnown, where the remaining known instances are retained, and RemoveKnown, where they are excluded from training.
Figure 2: Performance under two settings with different proportions of unknown knowledge in the same type and wiki test set.
As shown in Figure 2, the results across the four subplots are mutually corroborative, revealing a consistent pattern: the higher the proportion of unknown knowledge, the more severe the hallucination. In KeepKnown, performance on both the in-type test set and the OOD wiki test set degrades gradually at first, followed by a sharp decline as the amount of unknown knowledge increases. While RemoveKnown exhibits a more rapid degradation: within the target type, accuracy drops nearly linearly with the ratio of unknown samples; on the wiki set, performance declines sharply at low unknown counts and then plateaus. For the same replacement ratio, the key difference between KeepKnown and RemoveKnown lies in whether the type still contains known knowledge. We observe that this distinction has a substantial impact on model performance: RemoveKnown consistently underperforms KeepKnown. These observations suggest that sparse but fully unknown types are more disruptive than those containing a mixture of known and unknown knowledge, which differs from previous common understanding and poses new challenges for hallucination mitigation.

### v2

# Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis and Interpretation

@@ Abstract
Prior works have shown that fine-tuning on new knowledge can induce factual hallucinations in large language models (LLMs), leading to incorrect outputs when evaluated on previously known information. However, the specific manifestations of such hallucination and its underlying mechanisms remain insufficiently understood. Our work addresses this gap by designing a controlled dataset Biography-Reasoning, and conducting a fine-grained analysis across multiple knowledge types and two task types, including knowledge question answering (QA) and knowledge reasoning tasks. We find that hallucinations not only severely affect tasks involving newly introduced knowledge, but also propagate to other evaluation tasks. Moreover, when fine-tuning on a dataset in which a specific knowledge type consists entirely of new knowledge, LLMs exhibit elevated hallucination tendencies. This suggests that the degree of unfamiliarity within a particular knowledge type, rather than the overall proportion of new knowledge, is a stronger driver of hallucinations. Through interpretability analysis, we show that learning new knowledge weakens the model’s attention to key entities in the input question, leading to an over-reliance on surrounding context and a higher risk of hallucination. Conversely, reintroducing a small amount of known knowledge during the later stages of training restores attention to key entities and substantially mitigates hallucination behavior. Finally, we demonstrate that disrupted attention patterns can propagate across lexically similar contexts, facilitating the spread of hallucinations beyond the original task.
Understanding New-Knowledge-Induced Factual Hallucinations in LLMs:
Analysis and Interpretation
Renfei Dang1, Peng Hu1, Zhejian Lai1, Changjiang Gao1, Min Zhang2, Shujian Huang1 1 National Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China 2 Huawei Translation Services Center, Beijing, China {dangrf,hup,laizj,gaocj}@smail.nju.edu.cn, huangsj@nju.edu.cn, zhangmin186@huawei.com
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Large language models (LLMs) acquire rich factual knowledge during pre-training on massive text corpora (Petroni et al., 2019; Cohen et al., 2023), and are subsequently post-trained to follow human instructions and perform a wide range of downstream tasks (Ouyang et al., 2022; Wei et al., 2022).
@@
Figure 1: The impact of learning new knowledge on attention patterns and hallucination behavior. Training a model on new knowledge can induce factual hallucinations, whose severity is correlated with the model’s attention scores on key entities. Moreover, hallucinations are more likely to occur on test instances whose input contexts are lexically similar to the contexts of training tasks containing unknown knowledge. By injecting a small amount of known knowledge at the end of training, this issue can be effectively mitigated.
@@
However, during the Supervised Fine-Tuning (SFT) phase, models may encounter new knowledge not covered in pre-training. Prior researches (Ghosal et al., 2024; Lin et al., 2023; Ovadia et al., 2023; Gekhman et al., 2024; Sun et al., 2025) suggest that introducing new knowledge in the post-training phase increases the risk of factual hallucinations, where models generate fabricated yet plausible statements. This occurs because, when models learn new knowledge, they may erroneously generate related information in irrelevant contexts (Gekhman et al., 2024; Sun et al., 2025). These studies primarily focus on the effects within knowledge-intensive QA tasks during SFT, and we advance this line of research by investigating the fine-grained manifestations and underlying causes of hallucinations.
@@
To support this investigation, we construct a controlled experimental dataset Biography-Reasoning. The dataset is composed of biographical entities and their four attributes, which serve as four knowledge types. We further design twelve reasoning tasks using these knowledge. By controlling the proportion of known and unknown knowledge within different types and tasks in the training data, we systematically analyze the impact of learning new knowledge on hallucination risks.
@@
Our experiments reveal that training on unknown knowledge significantly elevates hallucination risks in the same task, while also inducing non-negligible hallucination effects on other out-of-domain test tasks. Importantly, we further find that when a knowledge type consists entirely of new knowledge, even a small amount of such data can markedly increase hallucination tendencies.
Through further interpretability analyses, we find that learning new knowledge significantly weakens the model’s attention to key entities in the question, thereby triggering factual hallucinations. In contrast, training on known knowledge strengthens the model’s attention to key entities. Motivated by this observation, we introduce a simple training method KnownPatch, which restores disrupted attention patterns by injecting a small amount of known knowledge during the later stages of training, and thus alleviates factual hallucinations. Finally, by constructing carefully designed variants of reasoning tasks, we demonstrate that lexical similarity (measured by token overlap between contexts), rather than semantic similarity of the contexts, is the primary driver of hallucination propagation across tasks. These findings are visually presented in Figure 1.
The main contributions of this paper are:
@@
  • Mechanism Interpretability: An analysis of attention mechanisms shows that learning new knowledge reduces attention to key question entities, causing hallucinations. In addition, lexically similar contexts facilitate the spread of these attention patterns, enabling cross-task hallucination effects.
@@ 2 Related Work
Existing studies have indicated that introducing new knowledge into LLMs may trigger hallucinations (Ghosal et al., 2024; Lin et al., 2023; Ovadia et al., 2023). Subsequent works have provided deeper analyses of this phenomenon (Gekhman et al., 2024; Kang et al., 2024; Sun et al., 2025). Gekhman et al. (2024) found that as the proportion of new knowledge in fine-tuning data increases, the model’s hallucination tendency intensifies. Kang et al. (2024) analyzed that when fine-tuned LLMs encounter unknown queries during testing, their responses imitate those associated with unknown examples in the fine-tuning data. From the perspective of token probabilities, Sun et al. (2025) show that after learning new knowledge, the generation probabilities of answer entity tokens increase significantly even in irrelevant contexts, suggesting that the model may over-generalize newly acquired knowledge and consequently produce hallucinations. However, previous studies focus mainly on closed-book QA settings with mixed knowledge types during training, while our controlled setup disentangles them to provide a more detailed analysis of new knowledge-induced hallucinations across types and tasks. Furthermore, we also investigate the underlying mechanisms of these phenomenon through an analysis of attention weights.
@@
Numerous studies are currently exploring ways to mitigate model hallucinations. A common approach involves providing additional relevant context to the model to reduce hallucinations during generation, such as through retrieval from knowledge bases or leveraging other large models to generate context (Shuster et al., 2021; Sun et al., 2022; Asai et al., 2024; Feng et al., 2023). Additionally, some research explicitly avoids hallucination risks by refusing to answer uncertain or unfamiliar questions (Yadkori et al., 2024; Zhu et al., 2025; Duwal, 2025). In another direction, many studies encourage the model to generate more known knowledge from pre-training, for example, by promoting factual outputs via reinforcement learning (Rafailov et al., 2023; Kang et al., 2024; Li and Ng, 2025; Gu et al., 2025) or by training only on known knowledge during supervised fine-tuning (Lin et al., 2024; Ghosal et al., 2024; Liu et al., 2024) to enhance the model. Our work builds on SFT with known knowledge approach, but rather than pursuing comprehensive filtering across all training data, KnownPatch only introduces a small number of known knowledge samples in the later stages of training, and alleviates the model’s tendency for hallucination.
@@ 3 Methodology of Analyzing Hallucinations
We aim to systematically investigate factual hallucinations in LLMs caused by learning different knowledge-related tasks. However, in real-world datasets, most factual knowledge may have already been seen by LLMs during pre-training, making it difficult to precisely control whether the knowledge being learned is new to the model. To address this limitation, we construct a synthetic dataset named Biography-Reasoning, which allows a controllable examination of hallucination behaviors under varying knowledge types and task types.
@@ 3.1 Biography-Reasoning Dataset
Following the data construction methodologies of Allen-Zhu and Li (2024); Zheng et al. (2025), we design the Biography-Reasoning dataset. The dataset centers on individuals as the key entities, with each person associated with four attributes: birth year, death year, major, and university. We refer to the same attribute of different individuals as a knowledge type.
Our dataset includes two types of knowledge-related tasks: knowledge QA and knowledge-based reasoning tasks. For knowledge QA tasks, questions are formulated by directly querying one of the attributes given the person’s name. Each task consists of questions on a single type, resulting in four QA tasks (e.g., Major_QA).
For knowledge-based reasoning tasks, we design three types of chain-of-thought-requiring reasoning tasks. Specifically, these include:
  • Single Reasoning: extracting one attribute from a single entity and performing a simple reasoning process;
  • Comparative Reasoning : extracting one attribute from each of two entities and performing comparative reasoning between them;
  • Novel Reasoning: extracting one attribute from a single entity and performing a newly defined reasoning task, such as mathematical or symbolic reasoning.
Table 1 presents examples of the constructed questions. The reasoning tasks are intentionally designed to be more complex than mere knowledge extraction as QA problems. Some of them require auxiliary knowledge (e.g., the major Dentistry belongs to the field Medicine), which the model is expected to contain. To further guarantee the model’s proficiency, we additionally collect and train on these auxiliary facts.
For each knowledge type we construct one QA and three reasoning tasks, leading to a total of 4 QA and 12 reasoning tasks per individual. Further dataset details can be found in Appendix A.
@@
QA Question: What major did Darreus Hsiao study?
@@
Single Reasoning Question: What field does Darreus Hsiao’s major belong to?
@@
Comparative Reasoning Question: Do Darreus Hsiao and Virgus Hong’s majors belong to the same field?
@@
Novel Reasoning Question: What is the sequence of odd-positioned letters in the first word of Darreus Hsiao’s major name?
@@ 3.2 Controlled Study Design
Since initially the model has no exposure to any knowledge of our synthetic dataset, we prepare the study by continue pre-training the model with a subset of the knowledge, which becomes known to the model; and keep another subset of the knowledge as unknown. By mixing the constructed questions from known and unknown knowledge in varying proportions, we are able to create situations where different proportion of newly introduced knowledge participates in training.
To evaluate how training leads to hallucinations, we reserve another subset of knowledge as test knowledge. The test knowledge are continue pre-trained together with the known knowledge during the preparation, but are kept away from further training. Therefore, the difference in performance on test set with and without unknown knowledge in the training data indicates the influence of factual hallucinations induced by training new knowledge. In addition, we use the real-world ENTITYQUESTIONS dataset (Sciavolino et al., 2021) derived from Wikidata (Vrandečić and Krötzsch, 2014) (denoted as Wiki) as an out-of-distribution (OOD) test set to provide a more robust evaluation.
@@ 3.3 Models and Setups
We conduct experiments primarily using the Qwen2.5-1.5B model (Team, 2024). As supplementary validation, we also perform key experiments on Llama3.2-1B (Grattafiori et al., 2024), Qwen3-8B-Base (Team, 2025) and Qwen2.5-32B (Team, 2024) to assess generalization across model scales and architectures, with their results provided in Appendix G.
@@
Following Allen-Zhu and Li (2024) and Gekhman et al. (2024), we adopt Exact Match (EM) as the metric for both knowledge QA tasks and reasoning tasks to assess the accuracy of the final answers. Given that all test knowledge are known to the model, and the training and testing formats are consistent, there are no cases where the answers are correct but incorrectly formatted. To ensure the generality of our conclusions, we report the standard deviation of accuracy where applicable.
@@ 4 Hallucination Analysis
Using the Biography-Reasoning dataset, we conduct a systematic study on factual hallucinations induced by learning different tasks containing various types of new knowledge through SFT. We additionally report the impact of learning new knowledge during CPT on hallucinations in Appendix E.
@@ 4.1 Knowledge QA
In this section, we analyze the impact of training on new knowledge in QA tasks. A baseline model is trained on samples constructed from the known knowledge of all four types. We then replace the knowledge of one entire type with unknown samples while keeping the other three types unchanged, resulting in four variant models. For each variant, we evaluate performance drop compared to the baseline model on three groups of QA test sets: (1) Same-Type QA (STQA): the QA test set whose knowledge type matches the type trained with unknown knowledge; (2) Different-Type QA (DTQA): the QA test sets whose knowledge types differ from the type trained with unknown knowledge; (3) Wiki: the real-world QA test set used as OOD evaluation.
@@
-54.26 ($\pm$ 3.30) -2.67 ($\pm$ 1.58) -17.72 ($\pm$ 6.56)
Table 2: Average performance degradation (%, mean with standard deviation) of four model variants. Detailed numerical results are reported in Appendix C.1.
Learning new knowledge induces factual hallucinations within the same type, with some spillover effects to other types. Table 2 presents the performance drop averaged across the four variant models. Training on unknown knowledge leads to substantial performance drops on the STQA test set, reducing the accuracy by more than half. We also observe cross-type degradation, as training on one type negatively impacts average performance on others, including the real-world Wiki test set containing OOD knowledge. This confirms that learning new knowledge can induce hallucinations even on unrelated knowledge. Notably, the performance drop on DTQA is smaller than on Wiki, as the former consists entirely of known data in the training set, which greatly mitigates the effect.
We further investigate how varying the proportion of unknown knowledge within a single type influences hallucination tendencies. Starting from the fully known-knowledge baseline, we progressively replace 5%, 10%, 20%, 50%, 80%, and 100% of the knowledge in one type with unknown knowledge, while still keeping the other three types entirely known. Two strategies are considered for handling the remaining known knowledge within the modified type: KeepKnown, where the remaining known instances are retained, and RemoveKnown, where they are excluded from training.
Figure 2: Performance under two settings with different proportions of unknown knowledge in the same type test set and Wiki test set.
As shown in Figure 2, the results across the four subplots are mutually corroborative, revealing a consistent pattern: the higher the proportion of unknown knowledge, the more severe the hallucination. In KeepKnown, performance on both the same-type test set and the OOD Wiki test set degrades gradually at first, followed by a sharp decline as the unknown knowledge dominants. In contrast, RemoveKnown exhibits a much faster degradation: even at low replacement ratios, the model already shows severe hallucination effects.
For the same replacement ratio, the key difference between KeepKnown and RemoveKnown lies in whether the knowledge type still contains any known instances. We observe that this distinction has a substantial impact on model performance, with RemoveKnown consistently underperforming KeepKnown. These observations suggest that sparse but fully unknown knowledge types are more disruptive than those containing a mixture of known and unknown knowledge, which differs from prior common understanding.
We additionally conduct the same set of experiments in this subsection using the real-world ENTITYQUESTIONS dataset. The results are reported in Appendix C.2, and all findings are highly consistent with those obtained on the synthetic data.